OpenStudy (anonymous):

Factor out the common factor of 3a. 9a^5/4-6a^9/2

OpenStudy (anonymous):

\[3a(\frac{3a^4}{4}-\frac{2a^8}{2})\]

OpenStudy (anonymous):

Is that what you want?

OpenStudy (anonymous):

\[3a^5(\frac{3}{4}-\frac{2a^4}{2})\]

OpenStudy (anonymous):

\[=3a^5(\frac{3}{4}-a^4)\]

OpenStudy (anonymous):

are the powers 5/4 and 9/2?

OpenStudy (anonymous):

yes they are the powers

OpenStudy (anonymous):

\[3^{5/4}(3-2a^{13/4})\]

OpenStudy (anonymous):

sorry,\[3a^{5/4}(3-2a^{13/4})\]

OpenStudy (anonymous):

bad day

OpenStudy (anonymous):

Are you factoring out 3a or 3a^5/4?

OpenStudy (anonymous):

If you are factoring 3a then the answer would be 3a(3a^1/4 - 2a^7/2)

OpenStudy (anonymous):

THanks beens.....I got another one if you dont mind

OpenStudy (anonymous):

no prob